If the economy is normal, Charleston Freight stock is expected to return 16.5 percent. If the economy falls into a recession, the stock's return is projected at a negative 11.6 percent. The probability of a normal economy is 80 percent while the probability of a recession is 20 percent. What is the variance of the returns on this stock? 


A. 0.010346 B. 0.012634 C. 0.013420 D. 0.013927 E. 0.014315
Answer #1

E(X) = sum of P(X) * X

= 0.8 * 0.165 + 0.2 * -0.116

= 0.1088

E(X^2) = sum of P(X) * X^2

=  0.8 * 0.165^2 + 0.2 * (-0.116)^2

= 0.0244712

Variance = E(X^2) - (E(x))^2

= 0.0244712 - 0.1088^2

= 0.012634

choose B)